# Emilia: An Extensive, Multilingual, and Diverse Speech Dataset for Large-Scale Speech Generation

### v2

@@ Abstract
Recently, speech generation models have made significant progress by using large-scale training data. However, the research community struggle to produce highly spontaneous and human-like speech due to the lack of large-scale, diverse, and spontaneous speech data. This paper presents Emilia, the first multilingual speech generation dataset from in-the-wild speech data, and Emilia-Pipe, the first open-source preprocessing pipeline designed to transform in-the-wild speech data into high-quality training data with annotations for speech generation. Emilia starts with over 101k hours of speech in six languages and features diverse speech with varied speaking styles. To facilitate the scale-up of Emilia, the open-source pipeline Emilia-Pipe can process one hour of raw speech data ready for model training in a few mins, which enables the research community to collaborate on large-scale speech generation research. Experimental results validate the effectiveness of Emilia. Demos are available at: https://emilia-dataset.github.io/Emilia-Demo-Page/.
@@ 1 Introduction
In the past few years, the research of speech generation has made significant advancements with the emergence of various generative models and the use of large-scale training data. The models such as Vall-E[1], SoundStorm [2], VoiceBox [3] and NaturalSpeech 3 [4] have considerably progressed in (zero-shot) speech generation by considerably scaling up both the datasets and model sizes, achieving high similarity, voice quality, and naturalness on academic datasets [5]. However, the generated speech still fails to generate speech akin to spontaneous human speech in the real world [4, 6].
@@
One of the significant reasons for this limitation is that current speech generation models are trained on speech datasets which have their root in audiobooks [7, 8]. Those datasets typically are characterized by formal reading styles. However, speech from real humans, especially in casual or conversational contexts, rarely adheres to such standardized patterns. Instead, it exhibits more diverse and spontaneous speaking styles, including breathing, pausing, repetitions, changes in speed, and varying emotions. Consequently, there is a pressing need for a new dataset that encompasses more diverse speech styles to advance the field towards generating more spontaneous and human-like speech.
@@
However, directly using in-the-wild speech data is not feasible due to variations in length and quality, frequent background noise, music, reverberation, the presence of multiple speakers within a single sample, and the lack of necessary annotations such as text transcriptions [9]. Training with such data may degrade the performance of speech generation models. While previous works [9, 10] propose automatic preprocessing pipelines to address these issues, they rely heavily on proprietary models, making their pipelines less accessible to the broader community. Additionally, the processing speed of these pipelines remains unknown. An ideal preprocessing pipeline for in-the-wild speech data should process quickly to handle large amounts of data efficiently, allowing for significant dataset scaling. Furthermore, the resulting datasets from these pipelines are limited to monolingual (Chinese-only) data and are relatively small in size (39 hours for [9], 12k hours for [10]).
In response to these issues, we present Emilia-Pipe, the first open-source preprocessing pipeline, which consists of six preprocessing steps: standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering. Emilia-Pipe can effectively transforms in-the-wild speech data into high-quality training data with annotations for speech generation. Additionally, Emilia-Pipe incorporates numerous engineering know-hows to improve robustness and efficiency. The resulting pipeline can process 2.50 hours of raw speech data in one minute using an independent server with eight NVIDIA RTX 4090 GPUs. It is also compatible with different languages.
Leveraging Emilia-Pipe, we construct the first multilingual speech generation dataset from in-the-wild speech data, Emilia. Table 1 compares Emilia with several existing speech generation datasets. The key advantages of the Emilia dataset are summarized as follows:
@@
  • Diverse. Unlike previous datasets, the Emilia dataset comprises mostly spontaneous speech, covering a wide range of speaking styles. This diversity is crucial for training high-quality models to generate spontaneous and human-like speech generation models.
  • Dynamic. The Emilia dataset features an automatic and efficient processing pipeline, allowing it to be easily extended in total duration and language coverage by adding user-specified source audios. To our best, this is the first dynamic speech dataset with an open-source data preprocessing pipeline, facilitating large-scale speech generation research for the community.
To validate the effectiveness of Emilia, we train two text-to-speech (TTS) models on the English subset of the Emilia dataset and compared them with their counterparts trained on the Multilingual LibriSpeech (MLS), a high quality dataset derived from audiobooks. Experimental results from both subjective and objective evaluations demonstrate that Emilia is effective for training high-quality, spontaneous, and human-like speech generation models. Additionally, models trained with the full Emilia dataset show promising performances in multilingual TTS.
The Emilia-Pipe and Emilia dataset, is now publicly available at https://github.com/open-mmlab/Amphion/tree/main/preprocessors/Emilia.
@@
2 The Emilia-Pipe Processing Pipeline
This section details Emilia’s data processing pipeline, Emilia-Pipe. As illustrated in Fig. 1, Emilia-Pipe includes six steps, i.e., Standardization, Source Separation, Speaker Diarization, Fine-grained Segmentation by VAD, ASR, and Filtering.
Fig. 1: An overview of the Emilia-Pipe processing pipeline.
@@ 2.1 Standardization
The raw speech data in the wild often vary in encoding formats, sampling rates, etc. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16-bit and adjust the target decibels relative to full scale to -20 dBFS. The actual gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. These steps ensure a consistent data format for further processing.
@@ 2.2 Source Separation
The raw speech data in the wild often contain background music, which negatively impacts the performance of speech generation models [16, 17]. To address this, we use the source separation technique to extract clean human vocals. Specifically, we utilize the open-source library, Ultimate Vocal Remover, [^1] https://github.com/Anjok07/ultimatevocalremovergui and its pre-trained model, UVR-MDX-Net Inst 3 [18]. [^2] https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models) This model achieves a high signal-to-distortion ratio of 11.15 for vocal separation on the Synth MVSep dataset. Using this model, we effectively separate human vocals from raw speech data for further processing.
@@ 2.3 Speaker Diarization
After extracting clean human vocals from the raw speech data, we apply the speaker diarization technique to partition the long-form speech data into multiple utterances based on the speaker. This process generates a series of utterances for each speech data, with each utterance containing only one speaker, ensuring compatibility with existing datasets for speech generation [7, 8, 12, 14]. To achieve this, we leverage the “pyannote/speaker-diarization-3.1” pipeline. [^3] https://github.com/pyannote/pyannote-audio This pipeline includes three core components: speaker segmentation, speaker embedding, and clustering, achieving state-of-the-art speaker diarization performance [19, 20]. The output of this pipeline is a list of temporal annotations indicating the start and end times of the single-speaker utterances.
@@ 2.4 Fine-grained Segmentation by VAD
Although the speaker diarization pipeline provides a coarse segmentation for the raw speech data, the resulting utterances may still be too long to fit into memory. To address this, we use a voice activity detection (VAD) model to further segment the utterances into smaller segments ranging from 3 to 30 seconds. This is achieved by concatenating consecutive chunks containing voice activity from the same speaker. We leverage the open-source library Silero-VAD. [^4] https://github.com/snakers4/silero-vad The pre-trained model provided in Silero-VAD achieves a ROC-AUC score of 0.99 on the LibriParty dataset, ensuring accurate detection of voice activity.
@@ 2.5 ASR
The absence of reliable text transcriptions impedes the direct use of the Emilia dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [21], which is a state-of-the-art multilingual ASR model capable of speech translation and language identification. To further enhance efficiency, we leverage the WhisperX [22], which builds on faster-whisper [^5] https://github.com/SYSTRAN/faster-whisper backend and the CTranslate2666https://github.com/OpenNMT/CTranslate2 inference engine, is up to four times faster than the official Whisper implementation for the same accuracy while using less memory. Additionally, we omit the original model’s VAD component by using the results from Sec. 2.4 to avoid redundant processing. We also develop batched inference for the faster-whisper backend to transcribe the speech data in parallel. These improvements, combined with the powerful Whisper model, allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.
@@ 2.6 Filtering
In real-world scenarios, some noise may not be effectively handled by source separation, the ASR step may introduce errors, and some raw speech data may be of low quality [9]. Therefore, to ensure the quality of the resulting dataset, we filter the data using the following criteria. [^7] Please note that the filtering criteria can be adjusted to fit the specific needs of different use cases. Firstly, we utilize the language identification results from the Whisper Model in Sec. 2.5. We discard the speech data that are not predicted to belong to our target languages (English, French, German, Chinese, Japanese, Korean) and have model language confidence lower than 80%. Secondly, we use DNSMOS P.835 OVRL score [23] to estimate the overall speech quality, preserving only those speech data with the score higher than 3.0. Finally, for each raw speech data, we compute the average character duration over its corresponding segments. We consider segments with an average phone duration outside 1.5 times the interquartile range (IQR) above the third quartile or below the first quartile as outliers and discard the speech data for these segments. After filtering, we obtain the resulting dataset, which are now ready for training the speech generation model.
@@ 3.1 Overview
Fig. 2: Duration statistics (hours) of the speech data by language.
Using Emilia-Pipe, we construct the Emilia dataset from a vast collection of speech data sourced from diverse video platforms and podcasts on the Internet, covering various content categories such as talk shows, interviews, debates, sports commentary, and audiobooks. This variety ensures the dataset captures a wide array of real human speaking styles.
After processing, the initial version of the Emilia dataset includes a total of 101,654 hours of multilingual speech data in six different languages: English, French, German, Chinese, Japanese, and Korean. Fig 2 provides the duration statistics for each language in the dataset.
@@ 3.2.1 Quality
To evaluate quality, we compared Emilia with existing datasets using DNSMOS P.835 OVRL scores. This non-intrusive speech quality metric reflects the overall quality of the speech data and is highly correlated with human ratings [23]. Table 3 presents the speech quality comparison between Emilia and several existing datasets.
Emilia achieves a DNSMOS P.835 OVRL score of 3.26, ranking third among all datasets. The results indicate that, despite being sourced from raw speech data in the wild, after preprocessing, the speech quality of the Emilia dataset is comparable to existing datasets sourced from studio recordings or audiobooks and outperforms all existing datasets sourced from in-the-wild speech data.
@@ 3.2.2 Diversity
For the semantic diversity analysis, we employ a pre-trained Sentence-BERT model [^9] https://github.com/UKPLab/sentence-transformers to generate text representations for the transcripts of each speech data. Consequently, each speech data is represented as a 768-dimensional vector based on its textual content, providing a comprehensive approximation of its semantic. Similar to the analysis above, we reduced the dimension of the semantic features to two. As shown in Fig. 3 (b), the scatter of textual features indicates that the Emilia dataset covers a wide range of textual content, validating the significant diversity in Emilia’s semantic coverage.
@@ 4 Experiments
In this section, we evaluate the effectiveness of the Emilia dataset in TTS applications through two experiments: the English-only experiment and the multilingual experiment. In the English-only experiment, we compare the performance of TTS models trained with the English subset of the Emilia dataset to those trained with the English subset of the MLS dataset. In the multilingual experiment, we train the models with the full Emilia dataset, which comprises 101,654 hours of multilingual speech data, and evaluate their multilingual TTS performance.
@@ 4.1.1 Baselines
In the experiments, we implement two TTS models as baselines: SoundStorm [2] and VoiceBox [16].
SoundStorm uses text to predict speech semantic tokens [25] in an autoregressive manner, and then generates acoustic tokens from a neural audio codec with bidirectional attention and confidence-based parallel decoding [2]. VoiceBox is a non-autoregressive speech synthesis model that adopts the non-autoregressive flow-matching method and the transformer for speech generation tasks. This model efficiently generates speech by learning the distribution of the mel-spectrogram conditioned on both text input and speech context.
@@ 4.2 Emilia English versus MLS English
The experiment evaluates the effectiveness of the proposed Emilia dataset by comparing the performance of models trained on the English subsets of both Emilia and the MLS dataset, a high-quality dataset derived from audiobooks. The total duration of the Emilia dataset is 46,000 hours, while the MLS dataset comprises 44,500 hours. The size of the datasets can be considered roughly equivalent. To thoroughly evaluate the models, we utilize the LibriSpeech-Test evaluation set, containing 1,200 speech samples in formal reading styles akin to those of MLS, and the Emilia-Test evaluation set, which includes 600 speech samples in diverse spontaneous speaking styles. Both test sets contain speech data that are unseen by the baselines.
Table 4 presents the results of both objective and subjective evaluations for the Emilia and MLS datasets on the LibriSpeech-Test and Emilia-Test. From these results, we observe that models trained on both the Emilia and MLS datasets demonstrate similar levels of speaker similarity (measured by SIM-O and SMOS) and intelligibility (measured by WER). This suggests that the Emilia dataset, despite being sourced from raw speech data in the wild, is as effective as high-quality datasets derived from audiobooks after processing with our proposed Emilia-Pipe. The VoiceBox model trained with both datasets demonstrates a similar level of naturalness. In contrast, the SoundStorm model shows significant improvement in FSD and CMOS on the Emilia-Test, which contains speech prompts in diverse spontaneous speaking styles. These results may indicate that autoregressive TTS models benefit more from speech with diverse speaking styles compared to non-autoregressive models.

### v3

@@ Abstract
Recent advancements in speech generation models have been significantly driven by the use of large-scale training data. However, producing highly spontaneous, human-like speech remains a challenge due to the scarcity of large, diverse, and spontaneous speech datasets. In response, we introduce Emilia, the first large-scale, multilingual, and diverse speech generation dataset. Emilia starts with over 101k hours of speech across six languages, covering a wide range of speaking styles to enable more natural and spontaneous speech generation. To facilitate the scale-up of Emilia, we also present Emilia-Pipe, the first open-source preprocessing pipeline designed to efficiently transform raw, in-the-wild speech data into high-quality training data with speech annotations. Experimental results demonstrate the effectiveness of both Emilia and Emilia-Pipe. Demos are available at: https://emilia-dataset.github.io/Emilia-Demo-Page/.
Index Terms—  Extensive Multilingual and Diverse Dataset, Large-scale Speech Generation
@@ 1 Introduction
In the past few years, the research of speech generation has made significant advancements with the emergence of various generative models [1, 2, 3, 4] and the use of large-scale training data [5, 6]. The models such as SoundStorm [1], VoiceBox [7] and NaturalSpeech 3 [3] have considerably progressed in (zero-shot) speech generation by considerably scaling up both the datasets and model sizes, achieving high similarity, voice quality, and naturalness on academic datasets [8]. However, the generated speech still fails to generate highly spontaneous and human-like speech in the real world [3, 9].
@@
One of the significant reasons for such limitation is that the current speech generation models are trained on speech datasets such as Libri-Light [5] and MLS [6], which have their root in audiobooks. Such datasets typically are characterized by formal reading styles. However, speech from real humans, especially in casual or conversational contexts, rarely adheres to such standardized patterns. Instead, it exhibits more diverse and spontaneous speaking styles, including breathing, pausing, repetitions, changes in speed, and varying emotions. Consequently, there is a pressing need for a new dataset that encompasses more diverse speech styles to advance the field towards generating more spontaneous and human-like speech.
@@
However, directly using in-the-wild speech data is not feasible due to variations in length and quality, frequent background noise, music, reverberation, the presence of multiple speakers within a single sample, and the lack of necessary annotations such as text transcriptions [10]. Training with such data may degrade the performance of speech generation models. While previous works [10, 11] propose automatic preprocessing pipelines to address such issues, they depend heavily on proprietary models, limiting their accessibility to the wider community. Additionally, the processing speed of these pipelines remains unknown. An ideal preprocessing pipeline for in-the-wild speech data should operate quickly to handle large volumes efficiently, enabling significant dataset scaling. Furthermore, the resulting datasets from these pipelines are limited to monolingual (Chinese-only) data and are relatively small in size (39 hours for [10], 12k hours for [11]).
In response, this paper introduces Emilia-Pipe, the first open-source preprocessing pipeline designed to efficiently convert in-the-wild speech data into high-quality, annotated training data for speech generation. Emilia-Pipe integrates several engineering optimizations to enhance both robustness and efficiency. The pipeline supports multiple languages and processes one hour of raw speech data in just a few minutes, enabling broader collaboration within the research community on large-scale speech generation. Leveraging Emilia-Pipe, we construct the first multilingual speech generation dataset from in-the-wild speech data, Emilia. Table 1 compares Emilia with several existing speech generation datasets. The key advantages of the Emilia dataset are summarized as follows:
@@
  • Diverse. The Emilia dataset comprises mostly spontaneous speech, covering a wide range of speaking styles. This diversity is crucial for training high-quality models to generate spontaneous and human-like speech generation models.
  • Dynamic. The Emilia dataset features an automatic and efficient preprocessing pipeline, Emilia-Pipe, which allows for easy scaling by incorporating readily available raw in-the-wild speech data or other user-specified source speech data.
To validate the effectiveness of Emilia, two text-to-speech (TTS) models are trained on the English subset of the Emilia dataset and compared with their counterparts trained on the MLS audiobook dataset. Experimental results from both subjective and objective evaluations demonstrate that Emilia is highly effective for training models capable of generating high-quality, spontaneous, and human-like speech. Furthermore, models trained on the full Emilia dataset show promising performance in multilingual TTS tasks.
The Emilia-Pipe and Emilia dataset, is now publicly available. [^1] GitHub: https://github.com/open-mmlab/Amphion/tree/main/preprocessors/Emilia; HuggingFace: https://huggingface.co/datasets/amphion/Emilia-Dataset.
@@
2 The Emilia-Pipe Preprocessing Pipeline
This section details Emilia’s data preprocessing pipeline, Emilia-Pipe. As illustrated in Fig. 1, Emilia-Pipe includes six steps, i.e., Standardization, Source Separation, Speaker Diarization, Fine-grained Segmentation by VAD, ASR, and Filtering.
Fig. 1: An overview of the Emilia-Pipe preprocessing pipeline.
@@ 2.1 Standardization
The raw speech data in the wild vary in encoding formats, sampling rates, etc. Therefore we convert all samples to WAV format, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16-bit and adjust the target decibel level to -20 dBFS. The volume adjustment is constrained between -3 and 3 dB to avoid distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. These steps ensure a consistent data format for further processing.
@@ 2.2 Source Separation
The raw speech data in the wild often contain background music/noise, which negatively impacts speech generation performance [2, 17]. To address this, we use the source separation technique to extract clean human vocals. Specifically, we utilize the Ultimate Vocal Remover, [^2] https://github.com/Anjok07/ultimatevocalremovergui and its pre-trained model, UVR-MDX-Net Inst 3333https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models, to separate human vocals from raw speech data for further processing.
@@ 2.3 Speaker Diarization
After extracting clean human vocals from the raw speech data, we apply the speaker diarization technique to partition the long-form speech data into multiple segments based on the speaker. This process generates a series of segments for each speech data, with each utterance containing only one speaker, ensuring compatibility with existing datasets for speech generation [5, 6, 13, 15]. To achieve this, we leverage the “pyannote/speaker-diarization-3.1” speaker diarization pipeline. [^4] https://github.com/pyannote/pyannote-audio This pipeline includes three core components: speaker segmentation, speaker embedding, and clustering, and achieves state-of-the-art speaker diarization performance [18]. The output of this pipeline is a list of temporal annotations indicating the start and end times of the single-speaker segments.
@@ 2.4 Fine-grained Segmentation by VAD
Although the speaker diarization pipeline provides a rough segmentation of the raw speech data, some segments may still be too long to fit into memory. To address this, we use the Silero-VAD model [^5] https://github.com/snakers4/silero-vad to further split any segment longer than 30 seconds into smaller VAD segments. We then concatenate consecutive VAD segments from the same speaker into appropriately sized utterances, ensuring each utterance is between 3 and 30 seconds in length.
@@ 2.5 ASR
The absence of text transcriptions limits the direct use of the Emilia dataset for TTS tasks. To address this, we apply ASR techniques to transcribe the segmented speech data. Balancing speed and accuracy, we utilize the state-of-the-art multilingual ASR model, Whisper-Medium. [^6] https://huggingface.co/openai/whisper-medium To further enhance efficiency, we use WhisperX [19], which is built on the faster-whisper [^7] https://github.com/SYSTRAN/faster-whisper backend and the CTranslate2888https://github.com/OpenNMT/CTranslate2 inference engine. This setup is up to four times faster than the official Whisper implementation while maintaining nearly the same accuracy. To avoid redundant processing, we bypass WhisperX’s VAD component by using the results from Sec. 2.4. Additionally, we implement batched inference with the faster-whisper backend to transcribe speech data in parallel. These optimizations significantly improves the efficiency of the pipeline.
@@ 2.6 Filtering
In real-world scenarios, some noise may not be fully addressed by source separation, the Whisper model may have hallucinations, and some raw speech data may be of low quality [10]. To ensure the quality of the resulting dataset, we apply the following filtering criteria. [^9] These criteria can be adjusted for the specific needs of different use cases. First, we use the language identification results from the Whisper model in Sec. 2.5 and discard any speech data not predicted to be in our target languages (English, French, German, Chinese, Japanese, Korean) or with model language confidence below 80%. Secondly, we use the DNSMOS P.835 OVRL score [20] to assess overall speech quality, retaining only speech data with a score higher than 3.0. Finally, for each raw speech sample, we calculate the average character duration across its corresponding segments. Segments with an average phone duration that falls outside 1.5 times the interquartile range (IQR) above the third quartile or below the first quartile are considered outliers, and the associated speech segments are discarded. After filtering, the resulting dataset is prepared for training the speech generation model.
@@ 3.1 Overview
Fig. 2: Duration statistics of the speech data by language.
Using Emilia-Pipe, we construct the Emilia dataset from a vast collection of speech data sourced from diverse video platforms and podcasts on the Internet, covering various content categories such as talk shows, interviews, debates, sports commentary, and audiobooks. This variety ensures the dataset captures a wide array of real human speaking styles. After processing, the initial version of the Emilia dataset includes a total of 101,654 hours of multilingual speech data in six different languages: English, French, German, Chinese, Japanese, and Korean. Fig 2 provides the duration statistics for each language in the dataset.
@@ 3.2.1 Quality
To evaluate quality, we compared Emilia with existing datasets using DNSMOS P.835 OVRL scores. This non-intrusive speech quality metric reflects the overall quality of the speech data and is highly correlated with human ratings [20]. Table 3 presents the speech quality comparison between Emilia and several existing datasets. Emilia achieves a DNSMOS P.835 OVRL score of 3.26, ranking third among all datasets. The results indicate that, despite being sourced from raw speech data in the wild, after preprocessing, the speech quality of the Emilia dataset is comparable to existing datasets sourced from studio recordings or audiobooks and outperforms all existing datasets sourced from in-the-wild speech data.
@@ 3.2.2 Diversity
For the semantic diversity analysis, we employ a pre-trained Sentence-BERT model [^11] https://github.com/UKPLab/sentence-transformers to generate text representations for the transcripts of each speech data. Consequently, each speech data is represented as a 768-dimensional vector based on its textual content, providing a comprehensive approximation of its semantic [22]. Similar to the analysis above, we reduced the dimension of the semantic features to two. As shown in Fig. 3 (b), the scatter of textual features indicates that the Emilia dataset covers a wide range of textual content, validating the significant diversity in Emilia’s semantic coverage.
@@ 4 Experiments
In this section, we evaluate the effectiveness of the Emilia dataset in TTS task through two experiments: the English-only experiment and the multilingual experiment. In the English-only experiment, we compare the performance of TTS models trained with the English subset of the Emilia dataset to their counterparts trained with the English subset of the MLS dataset. In the multilingual experiment, we train the models with the full Emilia dataset, which comprises 101,654 hours of speech data in six languages, and evaluate their multilingual TTS performance.
@@ 4.1.1 Baselines
We implement two TTS models as baselines: AR+SoundStorm [1] and VoiceBox [2]. AR+SoundStorm uses an autoregressive (AR) model to predict speech semantic tokens from text [23], followed by SoundStorm [1] to generate acoustic tokens in a non-autoregressive manner. VoiceBox is a fully non-autoregressive speech synthesis model that utilizes the flow-matching method and transformer architecture for speech generation, learning the distribution of the mel-spectrogram based on text input and speech context. For simplicity, we refer to AR+SoundStorm and VoiceBox as AR and NAR, respectively.
Both baselines are trained using four NVIDIA A100 GPUs for over 800k steps to ensure convergence. The models are optimized with the AdamW optimizer at a learning rate of 5e-5, including 5k warmup steps, followed by a cosine annealing learning rate schedule.
@@ 4.2 Emilia English versus MLS English
The experiment evaluates the effectiveness of the proposed Emilia dataset by comparing the performance of models trained on the English subsets of both Emilia and the MLS dataset, a high-quality dataset derived from audiobooks. The total duration of the Emilia dataset is 46k hours, while the MLS dataset comprises 44.5k hours. The size of the datasets can be considered roughly equivalent. To thoroughly evaluate the models, we utilize the LibriSpeech-Test evaluation set, containing 1,200 speech samples in formal reading styles akin to those of MLS, and the Emilia-Test evaluation set, which includes 600 speech samples in diverse spontaneous speaking styles. Both test sets contain speech data that are unseen by the baselines.
Table 4 presents the results of both objective and subjective evaluations for the Emilia and MLS datasets on the LibriSpeech-Test and Emilia-Test. From these results, we observe that models trained on both the Emilia and MLS datasets demonstrate similar levels of speaker similarity (measured by SIM-O and SMOS) and intelligibility (measured by WER). This suggests that the Emilia dataset, despite being sourced from raw speech data in the wild, is as effective as high-quality datasets derived from audiobooks after processing with our proposed Emilia-Pipe. The NAR model trained with both datasets demonstrates a similar level of naturalness. In contrast, the AR model shows significant improvement in FSD and CMOS on the Emilia-Test, which contains speech prompts in diverse spontaneous speaking styles. These results may indicate that AR TTS models benefit more from speech with diverse speaking styles compared to NAR models.